The following data show the age (week) and height (cm) of soybean 

plant. 

Age (week) 1 2 3 4 5 6 7

Height (cm) 5 13 16 23 33 38 40

Find the least square regression line of height on age. 


1
Expert's answer
2022-04-26T11:24:04-0400

Consider pairs of points "(x,y)", where "x" corresponds to the age and "y" corresponds to the height. The least squares regression line is the line: "\hat{y}=m{x}+b" that minimizes the following error: "E(m,b)=\sum_{i=1}^n(y_i-\hat{y})^2". After substitution "\hat{y}=mx+b" we receive: "E(m,b)=\sum_{i=1}^n(y_i-(mx_i+b))^2". After taking the derivatives with respect to "m" and "b" we get: "E_m=2\sum_{i=1}^nx_i((mx_i+b)-y_i)=2(m\sum_{i=1}^nx_i^2+b\sum_{i=1}^nx_i-\sum_{i=1}^nx_iy_i)" and "E_b=2\sum_{i=1}^n((mx_i+b)-y_i)=2(m\sum_{i=1}^nx_i+nb-\sum_{i=1}^ny_i)". Denote "\alpha=\sum_{i=1}^nx_i^2", "\beta=\sum_{i=1}^nx_iy_i", "\gamma=\sum_{i=1}^nx_i" and "\delta=\sum_{i=1}^ny_i". Necessary conditions of extrema take the form: "E_m=0,E_b=0". The latter is equivalent to: "m\alpha+b\gamma-\beta=0," "m\gamma+nb-\delta=0". From the first equation we get: "m=\frac{\beta-b\gamma}{\alpha}". We substitute it in the second equation and get: "\frac{\beta\gamma}{\alpha}-\frac{\gamma^2}{\alpha}b+nb-\delta=0"We receive: "b=\frac{\delta\alpha-\beta\gamma}{n\alpha-\gamma^2}", "m=\frac{n\beta-\delta\gamma}{n\alpha-\gamma^2}". Expressions for "m" and "b" as well as verifications of sufficient conditions of extrema can be also found in the respective literature about the least square regression line. Set "n=7", take values from the tables and get: "\alpha=140", "\beta=844", "\delta=168", "\gamma=28". We receive: "b=\frac{168\cdot140-844\cdot28}{7\cdot 140-(28)^2}=-\frac47", "m=\frac{7\cdot844-168\cdot28}{7\cdot 140-(28)^2}=\frac{43}{7}".

Answer: the least square regression line has the form: "\hat{y}=\frac{43}{7}x-\frac47". "x" corresponds to the age and "\hat{y}" corresponds to the height.